What are the lengths of the transverse and conjugate axes of the hyperbola given by the equation (y^2/9)-(x^2/10)=1?
Answer by lwsshak3(11628)

What are the lengths of the transverse and conjugate axes of the hyperbola given by the equation (y^2/9)-(x^2/10)=1?
This is an equation of a hyperbola with vertical transverse axis.
Its standard form: (y-k)^2/a^2-(x-h)^2/b^2=1, (h,k)=(x,y) coordinates of center.
For given equation:(y^2/9)-(x^2/10)=1
center: (0,0)
a^2=9
a=3
length of vertical transverse axis=2a=6
..
b^2=10
b=√10≈3.16
length of conjugate axis=2b=2√10≈6.32